Acidity (pH) of rain follows a normal distribution with a standard deviation of 3. In order to estimate the unknown population mean μ to within one unit with a confidence level of 98%, what would be the required sample size?

Select one:

a. 55

b. 49

c. 65

d. 99

e. 73
Answer #1

Solution :

Given that,

standard deviation =s =   =3

Margin of error = E = 1

At 98% confidence level the z is ,

= 1 - 98% = 1 - 0.98 = 0.02

/ 2 = 0.02/ 2 = 0.01

Z/2 = Z0.01 = 2.33 ( Using z table    )

sample size = n = [Z/2* / E] 2

n = ( 2.33* 3 / 1 )2

n =49

Sample size = n =49